OpenStudy (anonymous):

Solve -2(6-20n)=20(2n-3)

OpenStudy (saifoo.khan):

(no solutions exist)

OpenStudy (anonymous):

-12 + 40n = 40n - 60 -12 = -60 That is a contradiction. No solutions.

OpenStudy (anonymous):

solve for n that could be it or the foil method

OpenStudy (anonymous):

2(6-20n)=20(2n-3) 12-40n=40n-60 72=80n n=72/80 n=9/10